OpenStudy (anonymous):

\[\frac{ 6 }{ x ^{-5} }\] Convert this to the form \[{ ax ^{n} }\]

OpenStudy (anonymous):

\[\frac{1}{x^n}=x^{-n}\]

OpenStudy (anonymous):

i know about that rule, but what about the 6?

OpenStudy (anonymous):

\[\frac{6}{x^{-5}}=6\cdot\frac{1}{x^{-5}}\]

OpenStudy (anonymous):

Thanks very much!

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

@pgpilot326 another question: how does \[\frac{ 4 }{ 3\sqrt{x} } + \frac{ 5 }{ 2x ^{2} }\] = \[\frac{ 4 }{ 3}x ^{-\frac{ 1 }{ 2 }} + \frac{ 5 }{ 2}x ^{- 2 }\]